The sum of the first tow terms of a geometric sequence is 100 time the first term. What is the common ratio?

Answer by ikleyn(52790)

.
What do these words mean: "the first tow terms"?

Re-write and re-submit.

-------------------------------------------------------------------
Comment from student: The sum of the first two terms of a geometric sequence is 100 times the first term. What is the common ratio?
-------------------------------------------------------------------

OK.   So,  we have

a + a*q = 100*a,

where  a  is the first term of the geometric progression and  q  is the common ratio.

Divide both sides of this equation by  a.  You will get

1 + q = 100.

Hence,  q = 100 - 1 = 99.